OpenStudy (anonymous):

What is the value of z in the equation z + 7.8 = 16.4? A. 24.2 B. 7.6 C. 9.6 D. 8.6

OpenStudy (anonymous):

do opposite operation..

OpenStudy (anonymous):

\[z+7.8=16.4\]Subtract the 7.8 to the other side\[z=16.4-7.8\]Solve

OpenStudy (anonymous):

7.8-7.8=16.4-7.8

OpenStudy (anonymous):

Im sorry back

OpenStudy (anonymous):

8.6